OpenStudy (anonymous):

Medal Award Samantha purchased a dining room set for $2,910 using a 12-month deferred payment plan with an interest rate of 22.98%. She did not make any payments during the deferment period. What will Samantha’s monthly payment be if she must pay off the dining room set within two years after the deferment period? $121.25 $160.02 $191.33 $152.24

OpenStudy (amistre64):

compound the balance for the defered period, what do we have to work with?

OpenStudy (amistre64):

\[B_{12}=B_o~k^{12}~:~k=1+r/12\]

OpenStudy (amistre64):

using the value of B12, we can determine the Payments required for 24 months (2 years): \[[B]_{24}=B_{12}k^{24}-P\frac{1-k^{24}}{1-k}\]since we want the balance to be zero lets plug that in and solve for P \[0=B_{12}k^{24}-P\frac{1-k^{24}}{1-k}\] \[B_{12}k^{24}=P\frac{1-k^{24}}{1-k}\] \[B_{12}k^{24}\frac{1-k}{1-k^{24}}=P\] if we dont want to go thru the trouble of computing things twice, sub in for B12 = B0k^12 \[B_ok^{12}~k^{24}\frac{1-k}{1-k^{24}}=P\] \[B_ok^{36}~\frac{1-k}{1-k^{24}}=P~:~k=1+r/12\]

OpenStudy (anonymous):

I dont understand how to do it? im sorry im confused

OpenStudy (amistre64):

thats just the process; its the end formula i use to determine P http://www.wolframalpha.com/input/?i=2910k%5E%2836%29%281-k%29%2F%281-k%5E24%29%2C+k%3D1%2B.2298%2F12

OpenStudy (anonymous):

Thank you so much could you help me with another one?

OpenStudy (amistre64):

if it involve IRA stuff, then no ... cant seem to get those yet

OpenStudy (anonymous):

Alex purchased a bedroom set for $2,276 using a six-month deferred payment plan with an interest rate of 23.49%. What is the balance after the deferment period if payments of $112 are made each month? $1,604.00 $1,884.74 $2,276.00 $2,556.74

OpenStudy (anonymous):

i believe its b

OpenStudy (amistre64):

I use the same setup\[B_n=B_ok^n-P\frac{1-k^n}{1-k}~:~k=1+r/12\] since the defered period doesnt have a set payment; we can just set our own payment schedule ... they give it as P=112 \[B_6=2276~k^6-112\frac{1-k^6}{1-k}~:~k=1+.2349/12\] that comes to 1851 for me, so id say B is the safest bet http://www.wolframalpha.com/input/?i=2276k%5E6-112%281-k%5E6%29%2F%281-k%29%2Ck%3D1%2B.2349%2F12

OpenStudy (anonymous):

okay thanks :)

OpenStudy (amistre64):

good luck :)